OpenStudy (anonymous):

2 tan(3x)=2 over the interval [0,2pi)

OpenStudy (anonymous):

\[ \tan(3x)=1 \]Now we let \(u=3x\): \[ \tan(u) = 1 \]Solve this, and you get?

OpenStudy (anonymous):

 We get \(u=\pi/4\) and \(u=5\pi/4\).

OpenStudy (anonymous):

This means: \[ 3x = \pi/4,\quad 3x=5\pi/4 \]

OpenStudy (anonymous):

@jonbriley Do you follow?

OpenStudy (anonymous):